OpenStudy (anonymous):

A person starts his job with a certain monthly salary and earns a fixed increment every year. if this salary was rs 4500 after 4 years and rs 5400 after 10 years of service. find his initial salary and the annual increment.

OpenStudy (ankit042):

Try to convert this question as an arithmetic series {a+(n-1)d}

OpenStudy (anonymous):

couldnt do it! Its a linear equation thing

OpenStudy (ankit042):

how about this assume his initial salary as a and monthly increment as d now we have two unknowns and you can create two equations

OpenStudy (ankit042):

sorry yearly increment as d

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

salary is on monthly basis and increment on yearly

OpenStudy (ankit042):

hmm but increment will be for monthly salary so 4500 = a+3d and 5400 = a +9d

OpenStudy (anonymous):

how come?

OpenStudy (ankit042):

let us assume fixed increment is d rs pm(per month) and and his initial salary was a rs pm. Year Monthly salary First year a second year a+d third year a+2d fourth year a+3d